# TALES: A Taxonomy and Analysis of Cultural Representations in LLM-generated Stories

### v1

@@ Abstract.
Millions of users across the globe turn to AI chatbots for their creative needs, inviting widespread interest in understanding how such chatbots represent diverse cultures. At the same time, evaluating cultural representations in open-ended tasks remains challenging and underexplored. In this work, we present TALES, an evaluation of cultural misrepresentations in LLM-generated stories for diverse Indian cultural identities. First, we develop TALES-Tax, a taxonomy of cultural misrepresentations by collating insights from participants with lived experiences in India through focus groups (N=$9$) and individual surveys (N=$15$). Using TALES-Tax, we evaluate $6$ models through a large-scale annotation study spanning $2,925$ annotations from $108$ annotators with lived cultural experience from across $71$ regions in India and $14$ languages. Concerningly, we find that 8888% of the generated stories contain one or more cultural inaccuracies, and such errors are more prevalent in mid- and low-resourced languages and stories based in peri-urban regions in India. Lastly, we transform the annotations into TALES-QA, a standalone question bank to evaluate the cultural knowledge of foundational models. Through this evaluation, we surprisingly discover that models often possess the requisite cultural knowledge despite generating stories rife with cultural misrepresentations.
https://cultural-misrepresentations.github.io
copyright: acmlicensedjournalyear: 2025
@@ 1. Introduction
Figure 1. A broad overview of TALES: (RQ1) we identified categories of cultural misrepresentation through focus groups and surveys to develop TALES-Tax, (RQ2) conducted a large-scale annotation study to quantify the frequency of misrepresentation, and (RQ3) constructed TALES-QA from the annotated data to evaluate the cultural knowledge of models.:
@@ 3. RQ3:
To tackle RQ1, we examined cultural representations in LLM-generated stories through focus groups (N=$9$) and individual surveys (N=$15$), described in §3. We then qualitatively analyzed participants’ comments and thematically grouped them, resulting in TALES-Tax, a taxonomy of seven categories of cultural misrepresentations. The taxonomy is summarized in Table 1. While this taxonomy has been developed with a focus on India, its categories can be broadly used to evaluate cultural representation in generated text across other cultural contexts.
Next, we conducted a large-scale human evaluation of the prevalence of cultural misrepresentations in six popular LLMs, across English and $13$ other Indic languages (§4). For this, we employed $108$ annotators with lived cultural experience from $71$ regions and collected over $2,900$ span annotations of misrepresentations based on TALES-Tax across $540$ LLM-generated stories. Concerningly, we found that 8888% of the generated stories contained misrepresentations and they were more frequent in Indic languages and for peri-urban regions in India. Furthermore, based on participant ratings, we observed a noticeable scope for improvement in the relatability of the generated stories.
Finally, for RQ3, we converted the span annotations into TALES-QA, a set of standalone questions designed to test whether models simply lacked the knowledge that could have prevented the misrepresentations (§5). Surprisingly, despite prevalent misrepresentations in generated stories, the average accuracy across models on these questions was 76.976.9% in English and 59.859.8% in Indic languages. This finding suggests that most cultural misrepresentations cannot be explained merely by a lack of cultural knowledge; rather, such errors likely arise from the models’ inability to faithfully apply their stored cultural knowledge when producing open-ended, context-rich narratives.
@@
Overall, our work contributes to understanding cultural misrepresentations perpetuated by AI systems. Our evaluation is rooted in the rich perspectives of community members in both developing a taxonomy of cultural misrepresentations and conducting large-scale human evaluation of LLM-generated stories. Our findings underscore a gap in the competence of models in representing Indian cultures, especially in Indic languages and peri-urban regions. Moreover, we show that this is not due to a lack of knowledge pertaining to these identities. A broad overview of our methodology and analysis pipeline is shown in Figure 1. The data [^1] https://huggingface.co/datasets/Kirtibg/TALES-QA and code [^2] https://github.com/FLAIR-IISc/TALES is available publicly to support future research in this area.
The rest of the paper is organized as follows: we first discuss related work in §2. We then describe the methodology used to develop TALES-Tax in §3.1 along with a detailed description of the categories in §3.2 (RQ1). In §4, we describe our large-scale human evaluation (RQ2) with details about the annotation methodology (§4.1), evaluation methodology (§4.2), and findings (§4.3). Next, we evaluate LLM’s cultural knowledge in §5 (RQ3), with details about the creation of TALES-QA (§5.2), evaluation methodology (§5.2), and findings (§5.3). We conclude with a discussion on the implications, limitations, and future work.
@@ 2. Related Work
2.1. Evaluation of Cultural Competence of AI systems
@@
Cultural competence, also referred to as cultural alignment or cultural awareness, in contemporary literature is a growing area of research with broad goals to make AI systems “inclusive, adaptive, discerning, and nuanced” (Zhou et al., 2025a) to users from diverse sociocultural backgrounds. These goals of building AI systems that “work for everyone” are urgent, especially with the global adoption of AI technology (Mihalcea et al., 2025). We draw inspiration from the pyramid model of cultural competence proposed by Deardorff (2006). At the lowest level of the pyramid, AI systems must possess knowledge of diverse cultures. At higher levels, they must be able to apply this knowledge, be adaptable towards pluralistic cultural needs, and ultimately be able to effectively interact with or represent diverse cultures.
@@
Prior research has tackled evaluating cultural competence at various levels of this pyramid. Work that evaluated cultural knowledge in LLMs often involved creating datasets using various sources like expert documentation (Rao et al., 2024), social media platforms (Shi et al., 2024), knowledge resources (Zhao et al., 2024a; Zhou et al., 2025b), human-AI collaboration (Chiu et al., 2024), and crowdsourcing (Myung et al., 2024). Data resources have also been curated, targeted to the Indian context. For example, Seth et al. (2024) used participatory data curation to build a dataset for testing models’ knowledge of cultural entities from India. Similarly, the recently released SANSKRITI benchmark (Maji et al., 2025) is a collection of 2121K fact-based questions on India’s cultural diversity, compiled from $6$ online sources on topics such as tourism, cuisine, dance, and music. Our work adds to this through TALES-QA, a set of questions that test models’ cultural knowledge for diverse cultural identities. Although relatively smaller than SANSKRITI, our dataset is unique because it is grounded in deep cultural understanding of our expert annotators, who identified cultural misrepresentations in LLM-generated stories. Works have also evaluated models’ adherence to cultural values (AlKhamissi et al., 2024; Durmus et al., 2023) and adaptability towards cultural norms (Rao et al., 2024).
@@
Limited but growing attention is also being paid to cultural competence (or lack thereof) in generative applications. LLMs have been evaluated in the generative settings of open-ended question answering (Bhatt and Diaz, 2024), storytelling (Bhagat et al., 2025; Bhatt and Diaz, 2024), providing travel recommendations (Bhagat et al., 2025), and writing assistance (Agarwal et al., 2025). Our work adds to this body by focusing on cultural representation in storytelling. Past works that have considered this task (Bhagat et al., 2025; Bhatt and Diaz, 2024) have focused on global cultures and employed primarily quantitative evaluations. In contrast, we develop a fine-grained evaluation taxonomy through community engagement and conduct large-scale human evaluation focusing on cultural representations.
@@
Methodologically, culture has been operationalized through semantic proxies such as nationality (Bhatt and Diaz, 2024; Agarwal et al., 2025; Bhagat et al., 2025), food (Palta and Rudinger, 2023; Zhou et al., 2025b; Li et al., 2024b), names (Sandoval et al., 2023), values (Zhao et al., 2024a), or a combination (Seth et al., 2024). [^3] For more detailed breakdown see Adilazuarda et al. (2024) and Pawar et al. (2024) However, the limitations of such narrow operationalization are being increasingly recognized in the community (Zhou et al., 2025a). To overcome these limitations, some recent works have undertaken contextual qualitative analyses of cultural representation (Qadri et al., 2025, 2023). In this work, we combine the nuance of qualitative methods with the scale of quantitative analysis for evaluating cultural misrepresentations. Specifically, we first develop a taxonomy of cultural misrepresentations by conducting focus groups and individual surveys. We then base our larger-scale human evaluation of LLMs’ cultural misrepresentation across $14$ languages and identities on this taxonomy. This approach is closest to Bhatt et al. (2025), who evaluated alignment of LLMs to scientific communities by first eliciting requirements from experts through a qualitative study, followed by creating metrics for their criterion.
@@ 2.2. Story Generation and Evaluation
Recent work has found the growing role of LLMs in creative writing tasks such as story generation (Zhao et al., 2023). Beyond general-purpose models, specialized systems such as Weaver have been developed explicitly for content creation (Wang et al., 2024). To assess these advances, researchers have proposed diverse approaches to evaluate generated stories. Moving beyond lexical metrics (Xie et al., 2023), evaluations now consider dimensions such as creativity (Chakrabarty et al., 2024; Ismayilzada et al., 2024; Marco et al., 2024), coherence (Xu et al., 2018), story arcs (Tian et al., 2024), empathy and authenticity (Harel-Canada et al., 2024), and plot diversity (Xu et al., 2025).
@@
Despite these evaluation frameworks, existing works rarely account for whether generated stories represent diverse cultural knowledge. Recent work on measuring geographical bias in generated stories takes a quantitative approach by using a uniqueness score to capture cultural detail, showing that narratives about richer countries tend to be more distinctive than those about poorer countries (Bhagat et al., 2025). We build on this by incorporating user perspectives to evaluate the cultural representation in generated stories.
2.3. Categorization in Evaluation of AI Harms
@@
Empirical categorization has long been a cornerstone of research across disciplines. Organizing complex observations into coherent typologies creates structure for comparison, reveals blind spots, and enables cumulative knowledge building over time (Bailey, 1994). This tradition continues in contemporary AI research, where classification frameworks are used to make sense of the wide range of sociotechnical harms emerging from AI systems. For instance, scholars have developed taxonomies to describe sociotechnical risks (Shelby et al., 2023), user interaction patterns (Shelby et al., 2025), harms experienced by TGNB+ communities (Ungless et al., 2025), linguistic cues of anthropomorphism (DeVrio et al., 2025), forms of ableist hate (Heung et al., 2024), and risks related to text-to-image systems, privacy, and child safety (Bird et al., 2023; Lee et al., 2024; Basu et al., 2023; Mittal et al., 2024; Tsai et al., 2024; Mohseni et al., 2022).
@@
Within the growing literature on cultural competence in AI, researchers have similarly sought to survey and categorize how culture is represented, measured, and evaluated. Recent work has mapped proxies of culture, such as language, geography, and value dimensions—used in AI evaluations (Adilazuarda et al., 2024), and reviewed methodologies that aim to assess or improve cross-cultural awareness in AI systems (Pawar et al., 2024). While these efforts provide valuable overviews of how cultural factors are conceptualized, they remain largely descriptive and fall short of defining what it means for an AI system to be culturally representative in a specific context of use.
In particular, no existing framework articulates how to evaluate cultural competence in open-ended, creative applications, such as storytelling, where cultural appropriateness depends on nuanced understandings of local norms, symbols, and values. Our work addresses this gap by developing TALES-Tax, an empirically grounded taxonomy for evaluating the cultural appropriateness of generated stories, linking evaluation criteria to feedback from the communities whose cultures these systems aim to represent.
3. Identifying Categories of Cultural Misrepresentations (RQ1)
@@ 3.1. Methodology
To generate stories shown to participants during the focus groups and surveys, the research team curated a short list of four topics concerning festivals, wedding, childhood days, and daily lives. These topics cover aspects that vary socioculturally in India. For example, people celebrate different festivals across India, and the nature of celebration also varies considerably across cultures. Similarly, wedding rituals and daily lives can vary greatly based on region, religion, and other aspects of one’s cultural identity. To generate stories from an LLM, we created a prompt template that instructed an LLM to write a story about the given topic for a person with a specific cultural identity. For example, for someone from Chennai, the prompt would state: “Write a story about the childhood days of a person born and brought up in Chennai, India.” In line with prior work (Bhatt and Diaz, 2024), we intentionally used simple prompts to avoid additional confounding factors and focus our evaluation on cultural alignment in narratives rather than on the model’s ability to generate complex plots. For the focus groups and individual surveys, all stories were generated by GPT-4. We note that the findings of this part of the study are not a (quantitative) commentary on GPT-4’s ability to represent culture.
@@
Focus Groups (n=9)
24.6 ± 2.9
Male   5
Female  4
North  2  South  3
East  3  West  1
Surveys (n=15)
24.1 ± 2.8
Male   12
Female  3
North  5  South  7
East   1  West  2
@@
We conducted an hour-long group discussion with participants. One member of the research team acted as a facilitator and note-taker. Each group was presented with at least $3$ English stories each, which were generated to reflect the regions of its members. Stories were presented to the participants in a document that they could collaboratively comment on.
Participants first read each story independently and then engaged in a discussion, focusing on the representation of culture in the story. They were encouraged to identify inaccuracies, misrepresentations, and highlight parts of the story that felt nonsensical. To maintain engagement, the facilitator used a set of questions (see Appendix B) to guide participants to reflect on their cultural representation in the stories. All focus groups were conducted in English as it was the common language among all participants and the research team, and were recorded with participants’ consent.
@@
We also conducted surveys with individual participants to capture fine-grained assessments of misrepresentations in stories tailored to their cultural identities. Similar to the focus groups, each participant was presented with at least $3$ English stories constructed using the details of their cultural identities in a document. Each participant was asked to read the stories, highlight parts that they perceived as inaccurate or misrepresentative of their culture, and mention their reasoning in the document. A member of the research team was present when the participants participated in the survey to help them understand the exercise and clarify any doubts they had.
@@
To examine cultural misrepresentation in LLM-generated stories, we performed open-coding followed by reflexive thematic analyses (Braun and Clarke, 2022) of the discussion and comments from the focus groups and surveys following best practices. With the participants’ consent, we recorded the discussions during the focus group. We also collected the spans of stories that participants in focus groups and surveys highlighted as misrepresentations, and their comments describing their reasoning. Two authors and an external researcher, independently open-coded the spans and corresponding reasoning that the survey participants highlighted as misrepresentations. The first author additionally coded the transcripts of the focus group discussions. The authors met weekly during the coding process to group the $18$ initial codes into $7$ conceptual themes, resulting in TALES-Tax, our taxonomy of cultural misrepresentations. Following Braun and Clarke (2022), we did not calculate inter-rater agreement and instead relied on iterative discussion to reach consensus and finalize the categories of misrepresentations. Finally, one author coded all the data with the finalized categories to ensure that saturation had been reached. While this categorization reflects recurring themes in participants’ analyses of generated stories, we acknowledge that evaluating cultural representation is inherently subjective. Annotators drew on their own cultural experiences and perspectives when assessing the stories, and what one participant identified as inaccurate may be interpreted differently by another.
3.2. Findings: Categories of Cultural Misrepresentation (TALES-Tax)
Through the analyses of the participants’ comments in the focus group and surveys, seven categories of cultural misrepresentations emerged (see Table 1), which we now outline.

### v2

@@ Abstract.
Millions of users across the globe turn to AI chatbots for their creative needs, inviting widespread interest in understanding how they represent diverse cultures. However, evaluating cultural representations in open-ended tasks remains challenging and underexplored. In this work, we present TALES, an evaluation of cultural misrepresentations in LLM-generated stories for diverse Indian cultural identities. First, we develop TALES-Tax, a taxonomy of cultural misrepresentations by collating insights from participants with lived experiences in India through focus groups (N=$9$) and individual surveys (N=$15$). Using TALES-Tax, we evaluate $6$ models through a large-scale annotation study spanning $2,925$ annotations from $108$ annotators with lived experience and native language proficiency from across $71$ regions in India and $14$ languages. Concerningly, we find that 8888% of the generated stories contain misrepresentations, and such errors are more prevalent in mid- and low-resourced languages and stories based in peri-urban regions in India. We also transform the annotations into TALES-QA, a standalone question bank to evaluate the cultural knowledge of models.
Generative AI, storytelling, Human-AI interaction, Non-Western Cultures, India
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790519isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computingccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User studiesccs: Computing methodologies Natural language processing
@@ 1. Introduction
Figure 1. An overview of TALES: (RQ1) we identified categories of cultural misrepresentation through focus groups and surveys to develop our taxonomy, TALES-Tax, (RQ2) conducted a large-scale annotation study to quantify the frequency of misrepresentation, and (RQ3) constructed TALES-QA from the annotated data to evaluate the cultural knowledge of models.
Overview figure showing a three-stage research process arranged horizontally. The first stage depicts focus groups and surveys in which participants reviewed generated stories, and their feedback was analyzed to produce a taxonomy of seven types of cultural misrepresentation. The second stage shows a large-scale annotation process in which language experts used a span-annotation tool to mark and categorize misrepresentations in stories, with example spans highlighting incorrect geographic and cultural details. The third stage illustrates the transformation of annotated spans into a set of evaluation questions, with examples about traditional clothing and geographic locations used to assess models’ cultural knowledge.
@@ 3. RQ3:
To tackle RQ1, we examined cultural representations in LLM-generated stories through focus groups (N=$9$) and individual surveys (N=$15$), described in §3. We then qualitatively analyzed participants’ comments using reflexive thematic analysis, resulting in TALES-Tax, a taxonomy of seven categories of cultural misrepresentations. The taxonomy is summarized in Table 1. While this taxonomy has been developed with a focus on India, its categories can be broadly used to evaluate cultural representation in generated text across other cultural contexts.
Next, we conducted a large-scale human evaluation of the prevalence of cultural misrepresentations in six popular LLMs (§4). For this, we employed $108$ annotators with lived cultural experience and native language proficiency from $71$ regions and 14 languages across India. We collected over $2,900$ span annotations of misrepresentations based on TALES-Tax across $540$ LLM-generated stories. Concerningly, we found that 8888% of the generated stories contained misrepresentations and they were more frequent in Indic languages and for peri-urban regions in India. Furthermore, based on participant ratings, we observed a noticeable scope for improvement in the relatability of the generated stories.
Finally, for RQ3, we converted the span annotations into TALES-QA, a set of standalone questions designed to test whether models simply lacked the knowledge that could have prevented the misrepresentations (§5). Surprisingly, despite prevalent misrepresentations in generated stories, the average accuracy across models on these questions was 76.976.9% in English and 59.859.8% in Indic languages. This suggests that most cultural misrepresentations cannot be explained merely by a lack of cultural knowledge; rather, such errors likely arise from the models’ inability to faithfully apply stored cultural knowledge when producing open-ended, context-rich narratives.
@@
Overall,our work contributes to understanding cultural misrepresentations perpetuated by AI systems. Our evaluation is rooted in the rich perspectives of community members in both developing a taxonomy of cultural misrepresentations and conducting large-scale human evaluation of LLM-generated stories. Our findings underscore a gap in the competence of models in representing Indian cultures, especially in Indic languages and peri-urban regions. Moreover, we show that this is not due to a lack of knowledge pertaining to these identities. A broad overview of our methodology and analysis pipeline is shown in Figure 1. All the artifacts of the research, including the data and code, will be released publicly upon acceptance for utilization by future research.
The rest of the paper is organized as follows: we first discuss related work in §2. We then describe the methodology used to develop TALES-Tax in §3.1 along with a description of the categories in §3.2 (RQ1). In §4, we present our large-scale human evaluation (RQ2), including the annotation methodology (§4.1), evaluation methodology (§4.2), and findings (§4.3). Next, we evaluate LLM’s cultural knowledge in §5 (RQ3), detailing the creation of TALES-QA (§5.2), evaluation methodology (§5.2), and findings (§5.3). We conclude with discussion, limitations, and future work. All details about our work, including the code and data can be found through our website: cultural-misrepresentations.github.io.
@@ 2. Related Work
2.1. Evaluating Cultural Competence of AI
@@
Cultural competence, also referred to as cultural alignment or cultural awareness, in contemporary literature is a growing area of research with broad goals to make AI systems “inclusive, adaptive, discerning, and nuanced(Zhou et al., 2025a) to users from diverse sociocultural backgrounds. These goals of building AI systems that “work for everyone” are urgent, especially with the global adoption of AI technology (Mihalcea et al., 2025). We draw inspiration from the pyramid model of cultural competence proposed by Deardorff (2006). At the lowest level of the pyramid, AI systems must possess knowledge of diverse cultures. At higher levels, they must be able to apply this knowledge, be adaptable towards pluralistic cultural needs, and ultimately be able to effectively interact with or represent diverse cultures.
@@
Prior research has tackled evaluating cultural competence at various levels of this pyramid. Work has evaluated cultural knowledge in LLMs by creating datasets using various sources like expert documentation (Rao et al., 2024), social media platforms (Shi et al., 2024), knowledge resources (Zhao et al., 2024a; Zhou et al., 2025b), redteaming (Chiu et al., 2024), and crowdsourcing (Myung et al., 2024). Data resources have also been curated, targeted to the Indian context. For example, Seth et al. (2024) used participatory data curation to build ‘DOSA’, a dataset for testing models’ knowledge of cultural entities from India. Similarly, Maji et al. (2025) collect ‘SANSKRITI’, a dataset of 2121K fact-based questions on India’s culture, compiled from $6$ online sources on topics such as tourism, cuisine, dance, and music. Our work adds to this through a set of questions that test models’ cultural knowledge in multiple languages. Works have also evaluated models’ adherence to cultural values (AlKhamissi et al., 2024; Durmus et al., 2023) and adaptability towards cultural norms (Rao et al., 2024).
@@
Limited but growing attention is also being paid to cultural competence (or lack thereof) in generative applications. LLMs have been evaluated in the generative settings of storytelling (Bhagat et al., 2025; Bhatt and Diaz, 2024), providing travel recommendations (Bhagat et al., 2025), and writing assistance (Agarwal et al., 2025). Our work adds to this body by focusing on cultural representation in storytelling. Past works that have considered this task (Bhagat et al., 2025; Bhatt and Diaz, 2024) have focused on global cultures and employed primarily quantitative evaluations. In contrast, we develop a fine-grained evaluation taxonomy through community engagement and conduct large-scale human evaluation focusing on cultural representations.
@@
Methodologically, culture has been operationalized through semantic proxies such as nationality (Bhatt and Diaz, 2024; Agarwal et al., 2025; Bhagat et al., 2025), food (Palta and Rudinger, 2023; Zhou et al., 2025b; Li et al., 2024b), names (Sandoval et al., 2023), values (Zhao et al., 2024a), or a combination (Seth et al., 2024). [^1] For more detailed breakdown see Adilazuarda et al. (2024) and Pawar et al. (2024) However, the limitations of such narrow operationalization are being increasingly recognized in the community (Zhou et al., 2025a). To overcome these limitations, some recent works have undertaken contextual qualitative analyses of cultural representation (Qadri et al., 2025, 2023). In this work, we combine the nuance of qualitative methods with the scale of quantitative analysis for evaluating cultural misrepresentations. This approach is closest to Bhatt et al. (2025), who evaluated alignment of LLMs to scientific communities by first eliciting requirements from experts through a qualitative study, followed by creating metrics for their criterion.
@@ 2.2. Story Generation and Evaluation
Recent work has found the growing role of LLMs in creative writing tasks such as story generation (Zhao et al., 2023). Beyond general-purpose models, specialized systems like Weaver have been developed explicitly for content creation (Wang et al., 2024). To assess these advances, researchers have proposed diverse approaches to evaluate generated stories. Moving beyond lexical metrics (Xie et al., 2023), evaluations now consider dimensions such as creativity (Chakrabarty et al., 2024; Ismayilzada et al., 2024; Marco et al., 2024), coherence (Xu et al., 2018), story arcs (Tian et al., 2024), empathy and authenticity (Harel-Canada et al., 2024), and plot diversity (Xu et al., 2025).
@@
Despite these evaluation frameworks, existing works rarely account for how generated stories represent diverse cultural identities. Recent work on measuring geographical bias in generated stories takes a quantitative approach by using a uniqueness score to capture cultural detail, showing that narratives about richer countries tend to be more distinctive than those about poorer countries (Bhagat et al., 2025). We build on this by incorporating community perspectives to evaluate the cultural representation in generated stories.
2.3. Taxonomies in Evaluation of AI Harms
@@
Organizing complex observations into coherent taxonomies to create structure for comparisons, revealing blind spots, and enabling cumulative knowledge building over time has long been a cornerstone of research across disciplines (Bailey, 1994). This tradition continues in contemporary AI research, where taxonomies have been used to make sense of the wide range of sociotechnical harms emerging from AI systems. For instance, scholars have developed taxonomies to describe sociotechnical risks (Shelby et al., 2023), harms experienced by TGNB+ communities (Ungless et al., 2025), linguistic cues of anthropomorphism (DeVrio et al., 2025), forms of ableist hate (Heung et al., 2024), and risks related to text-to-image systems, privacy, and child safety (Bird et al., 2023; Lee et al., 2024; Basu et al., 2023; Mittal et al., 2024; Tsai et al., 2024; Mohseni et al., 2022).
@@
Within the growing literature on cultural competence in AI, researchers have similarly sought to survey and categorize how culture is represented, measured, and evaluated. Recent work has mapped proxies of culture, such as language, geography, and value dimensions—used in AI evaluations (Adilazuarda et al., 2024), and reviewed methodologies that aim to assess or improve cross-cultural awareness in AI systems (Pawar et al., 2024). While these efforts provide valuable overviews of how cultural factors are conceptualized, they remain largely descriptive and fall short of defining what it means for an AI system to be culturally representative in a specific context of use. In particular, no existing framework articulates how to evaluate cultural competence in open-ended, creative applications, such as storytelling, where cultural appropriateness depends on nuanced understandings of local norms, symbols, and values. Our work addresses this gap by developing TALES-Tax for evaluating the cultural appropriateness of generated stories, linking evaluation criteria to feedback from the communities whose cultures these systems aim to represent.
3. Developing a Taxonomy of Cultural Misrepresentations (RQ1)
@@ 3.1. Methodology
To generate stories shown to participants during the focus groups and surveys, the research team curated a short list of four topics concerning festivals, wedding, childhood days, and daily lives. These topics cover aspects that vary socioculturally in India. For example, people celebrate different festivals across India, and the nature of celebration also varies considerably across cultures. Similarly, wedding rituals and daily lives can vary greatly based on region, religion, and other aspects of one’s cultural identity. To generate stories from an LLM, we created a prompt template that instructed an LLM to write a story about the given topic for a person with a specific cultural identity. For example, for someone from Chennai, the prompt would state: “Write a story about the childhood days of a person born and brought up in Chennai, India.” For the focus groups and individual surveys, all stories were generated by GPT-4. We note that the findings of this part of the study are not a quantitative commentary on GPT-4’s ability to represent culture.
@@
Focus Groups (n=9) 24.6 ± 2.9 Male   5 Female  4 North  2  South  3 East  3  West  1
Surveys (n=15) 24.1 ± 2.8 Male   12 Female  3 North  5  South  7 East   1  West  2
@@
We conducted an hour-long group discussion with participants. One member of the research team acted as a facilitator and note-taker. Each group was presented with at least $3$ English stories each, which were generated to reflect the regions of its members. Stories were presented to the participants in a document that they could collaboratively comment on. Participants first read each story independently and then engaged in a discussion, focusing on the representation of culture in the story. They were encouraged to identify inaccuracies, misrepresentations, and highlight parts of the story that felt nonsensical. To maintain engagement, the facilitator used a set of questions (see Appendix B) to guide participants’ to reflections. All focus groups were conducted in English as it was the common language among all participants and the research team.
@@
We also conducted surveys with individual participants to capture fine-grained assessments of misrepresentations in stories tailored to their cultural identities. Similar to the focus groups, each participant was presented with at least $3$ English stories constructed using the details of their cultural identities in a document. Each participant was asked to read the stories, highlight parts that they perceived as inaccurate or misrepresentative of their culture, and mention their reasoning. A member of the research team was present when the participants participated in the survey to help them understand the exercise and clarify any doubts.
@@
To examine cultural misrepresentation in generated stories, we performed open-coding followed by reflexive thematic analyses (Braun and Clarke, 2022) of the discussion and comments from the focus groups and surveys following best practices. With the participants’ consent, we recorded the discussions during the focus group. We also collected the spans of stories that participants in focus groups and surveys highlighted as misrepresentations, and their comments describing their reasoning. Two authors and an external researcher, independently open-coded the spans and corresponding reasoning that the survey participants highlighted as misrepresentations. The first author additionally coded the transcripts of the focus group discussions. The authors met weekly during the coding process to group the $18$ initial codes into $7$ conceptual themes, resulting in TALES-Tax, our taxonomy of cultural misrepresentations. While these themes reflects our participants’ analyses of generated stories, we note that evaluating cultural representation is inherently subjective. Participants drew on their own cultural experiences and perspectives when assessing the stories, and what one participant identified as inaccurate may be interpreted differently by another. Thus, following Braun and Clarke (2022), we did not calculate inter-rater agreement, but resolved disagreement or subjective interpretations through iterative discussions till agreement among the research team was reached. Finally, one author used the 7 categories to code all of the data to ensure that no additional misrepresentations emerged, ensuring saturation.
3.2. Findings: Taxonomy of Cultural Misrepresentations (TALES-Tax)
Through thematic analyses of the participants’ comments in the focus group and surveys, seven categories of cultural misrepresentations emerged (see Table 1).